40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9659     Entry time: Wed Feb 19 22:47:26 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: ALS locked using LSC model, Common & Diff transitioned to IR transmission signals 

[Jenne, Koji, Manasa, EricQ]

Today we successfully locked the ALS using the LSC system, with filters that are good for both the IR PDH and the ALS locking.  We tried PRFPMI, but were unable to hold PRMI lock while the arms were held with ALS.  We combined the ALS signals into common and differential signals, and successfully transitioned over to a combined set of 1/sqrt(TRANS) signals for the common mode part of the lock (differential stayed with ALS). 


Locking the ALS using filters in the LSC system that are also good for IR PDH

The biggest difference between the ALS and LSC filters were the ones used for lock aquisition. At Koji's suggestion, I made FM5 of the LSC servos (for X and Y arms) the filter needed for ALS locking.  Then, I made FM4 into a combination of old LSC FM4 and FM5, as well as an inverse of the new FM5, so that when both FM4 and FM5 are engaged, the servo shape is the same as the old LSC.  I left the other LSC filters where they were.  I replaced the FM1 +6dB with the combined integrators (really, just gentle DC boosts) for the ALS, since we were never using this +6dB filter module.  The LSC resonant gain filter for the bounce mode also included a resgain for 18.5 Hz.  I don't know what that was for, and it was eating into phase that I needed, so I removed it.

The other filter that changed significantly was the Boost filter.  The ALS system had been using more DC gain than the LSC had.  However, the current ALS boost filter (in FM10 of the old ALS servos) was eating too much phase near my UGF.  So, I scooted the whole boost filter to lower frequencies, to give myself some extra phase margin.  The boost was set to "zero history", "zero crossing", with 0.01 tolerance and an 8 second timeout.  Setting it to zero crossing with a low tolerance, rather than just ramping it on, was the key to engaging the boost.

ALS_newVSoldBoosts.pdf

I had to be so careful about phase margin, since I lost ~15 degrees of phase at 200 Hz from the lag of going through the RFM network.  This was pretty frustrating, but I don't have a better plan yet, save moving the c1als model and ADC to the SUS machine, which has Dolphin access to the LSC.  I may back off my safety margin, and give myself some gain in the boost back at 10Hz, since we are now seeing too much noise at 10Hz in the closed-loop spectra.  I also "cheated" and lowered my UGF from the ~150Hz it used to be in the ALS model, to 100Hz, where I was closer to the top of the new phase bubble.

With the new filter situation, I was able to lock the Xarm (the one I was using for design work) with both IR and ALS.  To lock IR, the "restore" script still works. For the ALS, we should put in a separate "restore" script into the IFO_CONFIGURE screen. 

The ALS locking procedure is as follows:

* Prepare ALS and green locking.  Green locked to 00 mode, alignment all nice, etc, etc.  Beatnote within 100MHz on spectrum analyzer.  If doing both arms, try to get beatnotes on opposite sides of PSL, to keep crossbeatnotes at higher frequencies, and out of the way.

* Turn on Watch script.

* Set LSC parameters (this is where a new restore script will come in handy): 

       * Zeros in RFPD columns of input matrix (i.e. POX and POY).

       * Ones in AUX input matrix elements.

       * Zeros in power normalization matrix rows for arms.

       * All FM triggers for arms set to "Man" for manual.

       * Override main trigger, so that signals are always going through to the servo.

       * Only FM5 engaged in arm servo.

       * Gain of servo set to zero, output on, then engage main LSC master switch.  ETM output on.

* Clear history in phase tracker.

* Check sign of gain using + or - 0.1 in the servo.  You'll know if you got it wrong (the ETM will be kicked, and the beatnote will fly around).  If you didn't get it wrong, you probably got it right.

* Increase gain to about 12 (with correct sign).

* Engage FM1 (gentle DC boost), FM6,7,8 (resonant gains for stack, bounce, roll)

* Wait a few seconds for filters to settle, then engage FM9 (boost).

* Run find IR resonance script.

* Move off resonance by ~36 counts (12 times the +3 script).  This number comes from trying to be completely off the IR resonance, even when the PRMI was locked.

* Do whatever locking (ex. PRMI) you set out to do.


 PRFPMI attempt

After locking both arms with ALS using the LSC system, we attempted to lock the PRMI.  We were able to lock PRMI on REFL55 I&Q, REFL33 I&Q, and REFL55 I&AS55Q before the arms were locked, so we were hoping that we wouldn't have too much trouble.

We found the IR resonance for both arms, then moved off resonance.  Then, restored the PRM.  For REFL55, Koji coarsely turned the REFL 55 demod phase from 16 degrees to 87, while we were locked on the carrier.  After this, I stepped farther and farther from the IR resonance, since at first I found that our transmitted powers were something like 4, rather than almost zero, so the demod phase may not be totally correct.  

We were having trouble, so we locked the PRMI on carrier using REFL55 I and AS55 Q, with 1's in both elements in the input matrix.  MICH gain was about -10, PRCL +0.010.  We used this time to tweak up the alignment of the PRMI.  At some point, Koji tweaked the REFL33 demod phase from 124 to 134 degrees.  Then we switched back to sideband locking.  After some trials with REFL55 I&Q, and REFL55/AS55, we went to REFL33 I&Q.  REFL33I->PRCL was 1.556 in the input matrix, and REFL33Q->MICH was -0.487.  No other elements in the input matrix.  MICH gain was reduced to -6, PRCL gain to -0.020.  MICH FMs 3,6,9 triggered, PRCL FMs 2,3,6,8,9 triggered.  We were able to keep short locks on the order of ~10 seconds, but not longer. We played with every parameter we could think of (alignment being good is one of the most important!), but were not able to keep better lock.  The POP spot is moving around a lot, so the PRCL ASC needs to be examined, hopefully tomorrow.

We started losing the Xarm lock fairly regularly, I'm not sure why, but the Yarm was locked for almost 2 hours straight, held off resonance with ALS!


 ALS Common and Differential, transition to IR control

We set PRMI aside for the rest of the night, and looked at using ALS to control the arms in common and differential modes. 

Regular ALS locking procedures were used (see above), with the exception of the AUX input matrix:

  1/sqrt(TRX) 1/sqrt(TRY) ALSX ALSY
XARM (common) 0 0 +1 -1
YARM (differential) 0 0 +1 +1

 Since the beatnotes were on opposite sides of the PSL frequency, the common and differential modes look opposite of what you'd expect. 

We then used the regular find IR resonance scripts running simultaneously, which worked really well to find both arms' IR resonance points.

I put a 1 count offset in the Xarm servo (which was our proxy for common mode), although in retrospect this should have been +0.5 in ALSX, and -0.5 in ALSY, so that our signals going through the input matrix were at their zero crossings.  Anyhow, this offset put us at about half fringe on both arms (transmissions were about 0.6). 

Koji set the offsets in the 1/sqrt(trans) filter banks before the input matrix so that they would have zero crossings at this point (avg the IN1, put negative of that value into the offset). 

We then stepped the input matrix values until our common mode (Xarm) row was:

  1/sqrt(TRX) 1/sqrt(TRY) ALSX ALSY
XARM (common) -0.7 -0.7 0 0

We left the differential (YARM) row alone, so that the ALS system would still be controlling the differential degree of freedom.  The values and sign for the 1/sqrt(trans) signals came from a transfer function of dividing the spectra of each error signal and noting the relative gain and sign.

After we swapped the error signals, we realized that we had to remove the offset from the XARM servo, which is why we should have put the offsets elsewhere in the first place.

Then, Koji took a spectrum, which is attached to this entry.  We note that the ALS signals are strongly correlated, and mostly common. 


To Do List

Going forward, we need to figure out what is going on with the PRMI, and why we're having trouble keeping lock.

We need to redo the PRCL ASC servo, with the anti-oplev trick that Rana mentioned a week or two ago.

We need to investigate the degeneracy of REFL165, now that Q's simulation doesn't justify / explain it. 

Attachment 1: common_diff_ALS.pdf  108 kB  | Hide | Hide all
common_diff_ALS.pdf
ELOG V3.1.3-